OpenStudy (anonymous):

Solve by the multiplication with add or subtract method 5x + 2y = -8 3x + 7y = 1

OpenStudy (texaschic101):

5x + 2y = -8 ---->(-3) 3x + 7y = 1----->(5) ----------------- -15x - 6y = 24 (result of multiplying by -3) 15x + 35y = 5 (result of multiplying by 5) ----------------add 29y = 29 y = 29/29 y = 1 now sub 1 in for y in either equation to find x 5x + 2y = -8 5x + 2(1) = -8 5x + 2 = -8 5x = -8 - 2 5x = -10 x = -10/5 x = -2 check.. 3x + 7y = 1 3(-2) + 7(1) = 1 -6 + 7 = 1 1 = 1 (correct) x = -2 and y = 1 or (-2,1)

OpenStudy (anonymous):

Can you help me with 3x - 4y = 45 7x + 6y = 13

OpenStudy (texaschic101):

3x - 4y = 45 --->(3) 7x + 6y = 13 --->(2) ------------------ 9x - 12y = 135 (result of multiplying by 3) 14x + 12y = 26 (result of multiplying by 2) ------------------add 23x = 161 x = 161/23 x = 7 now sub 7 in for x in either equation to find y 3x - 4y = 45 3(7) - 4y = 45 21 - 4y = 45 -4y = 45 - 21 -4y = 24 y = 24/-4 y = - 6 check... 7x + 6y = 13 7(7) + 6(-6) = 13 49 - 36 = 13 13 = 13 (correct) x = 7 and y = -6 or (7,-6)